OpenStudy (anonymous):

is f(x) = -3x + 8 odd even or neither function

jhonyy9 (jhonyy9):

add values to x and make the calcule check what number will get when x=0 than f(0)= -3*0 +8 = 0 +8 =8 so this is even continue and see the result

OpenStudy (anonymous):

is that the definition of odd or even function? A check of the value of the function at f(0)? I have been confused on that before...

hartnn (hartnn):

Replace x in f(x) by -x to find f(-x) if f(x)=f(-x) then function is even if f(x)=-f(-x) then function is odd else its neither

OpenStudy (anonymous):

@hartnn thanks!

OpenStudy (anonymous):

so a function like f(x)= x^3 is odd, and f(x) = x^2 would be even? (not strictly because of the exponents being odd or even, although in this simplistic example, that is also true...) Thanks again... glad to have restored some lost knowledge!!